OpenStudy (toxicsugar22):

1.put the equation 2y-3x=5 into slope intercept form 2.use the point slope formula to write an eqaution for the line that is perpendicular to the graph of 2y-3x=5 and passes through the point (1,4)

OpenStudy (anonymous):

Slope-int form: y = mx +b 2y-3x = 5 2y = 3x +5 y = 3x/2+5/2 2. m = 3/2 Perpendicular slope = -2/3 y - 4 = (-2/3)(x-1)

OpenStudy (toxicsugar22):

thanks